# Sequential Diversification with Provable Guarantees

### v2

@@ 15            $\pi(2{\tau}-1)\leftarrow v_{2{\tau}-1}$, $\pi(2\tau)\leftarrow u_{2{\tau}-1}$
16      if $d(v_{2i-1},\pi(2i+1))\geq d(u_{2i-1},\pi(2i+1))$ then
@@ 6.2. The greedy-matching algorithm
In this section, we examine the case where $p$ is large, specifically when $\lim_{n\rightarrow\infty}p=1$ and $\lim_{n\rightarrow\infty}p^{n}=0$. This suggests that as the number of items increases, the user accepts more items, but eventually quits before examining all of them. For instance, setting $p=1-\frac{1}{\log\log n}$ implies that, in expectation, the user accepts $\log\log n$ items before quitting. For simplicity, we assume $p=1-\frac{1}{t_{n}}$, where $t_{n}$ is an increasing function of $n$.
In this scenario, applying $\textsc{B}\tau\textsc{I}$ algorithm to find best-$\tau$ items becomes computationally infeasible. As Theorem 1 shows, achieving a constant approximation requires setting $\tau=\Omega(t_{n})$. To address this, we introduce a polynomial-time greedy-matching (GM) algorithm.
@@
The next step is to extend $M$ to a Hamiltonian path. Notice that the (undirected) edge $(u_{2j-1},v_{2j-1})$ is always the $(2j-1)$-th edge on the path, and the algorithm needs to decide which item comes first in the path. To make this decision, the algorithm utilizes the triangle inequality.
@@ 7. Non-uniform continuation probabilities
In this section, we discuss a more general case when the continuation probabilities $p_{i}$ are non-uniform. We assume $p_{i}\in[a,b]$, for all $i\in U$, where $10<a\leq b<$ 1. We start by adapting the $\textsc{B}\tau\textsc{I}$ algorithm, introduced in Section 6, to this setting. Given the computational complexity of the $\textsc{B}\tau\textsc{I}$ algorithm, we also propose an efficient greedy algorithm, whose performance guarantee is an implication of the $\textsc{B}\tau\textsc{I}$ algorithm by setting $\tau=2$.
@@ 7.1. The best-$\tau$ items algorithm
Recall that the ordered Hamiltonian-path is defined as $\mathcal{H}(O)=\sum_{i=1}^{n-1}W_{O_{i}}d(\pi(i),\pi(i+1))$, where $W_{O_{i}}=\sum_{j=i+1}^{n}p_{O_{j}}$. When the continuation probabilities $p_{i}$ are uniformly equal to $p$, we observe that the value of $W_{O_{i}}$ is dominated by $\nicefrac{{p^{i+1}}}{{(1-p)}}$, hereby, we can use $\nicefrac{{p^{i+1}}}{{(1-p)}}$ to approximate $W_{O_{i}}$. However, when $p_{i}$ are not uniform, it is hard to find a simple and closed-form equation to approximate $W_{O_{i}}$, which is convenient for our analysis.
@@
~$\tilde{\mathcal{H}}(O_{\tau})=\sum\nolimits_{i=1}^{\tau-1}\sum\nolimits_{j=i+1}^{\tau}p_{O_{j}}d(\pi(i),\pi(i+1)),$
and we find the best-$\tau$ items by maximizing ~$\tilde{\mathcal{H}}(O_{\tau})$. We redefine
@@
The $\textsc{B}\tau\textsc{I}$ algorithm (Algorithm 1) has the following quality guarantee.
@@
Let $p_{i}\in[a,b]$ for all $i\in U$, where $10<a\leq b<$ 1. Let $\tau$ be an integer such that $2\leq\tau\leq n$. The $\textsc{B}\tau\textsc{I}$ algorithm provides a $\frac{a^{2}(1-b)(1-b^{\tau-1})}{a^{2}+(\tau-1)b^{\tau+1}}$-approximation guarantee for MaxOHP.
@@ 7.2. The greedy algorithm
Since the $\textsc{B}\tau\textsc{I}$ algorithm is impractical when $\tau$ is large, we propose a greedy algorithm, presented in Algorithm 3. Its approximation ratio is obtained by setting $\tau=2$ in Theorem 1.
@@ 8. Experimental evaluation
Our implementation and the datasets are publicly available. [^1] https://github.com/HongLWang/Sequential-diversification-with-provable-guarantees
@@ 8.1. Datasets
We evaluate our methods on seven public datasets: five commonly used in recommender systems and two, LETOR and LTRC, from the information retrieval field. We refer to these datasets as recommendation datasets and information retrieval datasets, respectively. Table 1 provides a summary of dataset statistics. Due to space constraints, the dataset URLs are provided in the appendix.
@@
Table 1. Dataset statistics of the $5$ recommendation datasets and $2$ information retrieval datasets. $\mathcal{U}$ and $\mathcal{I}$ represent the set of users (or queries) and items (or documents), respectively. $\#\mathcal{R}$ denotes the number of ratings (or relevance scores), $\mathrm{avg}(\mathcal{R})$ denotes the average number of ratings per user (or documents per query), and $\mathrm{avg}(dist)$ denotes the average item-item (or document-document) distance.
Datasets $|\mathcal{U}|$ $|\mathcal{I}|$$\#\mathcal{R}$ $\mathrm{avg}(\mathcal{R})$ $\mathrm{avg}(dist)$
@@
[…]
Datasets Random EXPLORE DUM MSD MMR DPP B2I B3I B4I
Coat 0.646 $\pm$ 0.177 1.276 $\pm$ 0.374 1.214 $\pm$ 0.348 1.211 $\pm$ 0.349 1.282 $\pm$ 0.377 1.284 $\pm$ 0.377 1.289 $\pm$ 0.376 1.184 $\pm$ 0.368* 1.184 $\pm$ 0.367*
KuaiRec 0.383$\pm$ 0.066 0.681 $\pm$ 0.082 0.411 $\pm$ 0.086 0.704 $\pm$ 0.057 0.725 $\pm$ 0.080 0.711 $\pm$ 0.073 0.778 $\pm$ 0.075 0.776 $\pm$ 0.072* 0.782 $\pm$ 0.073*
Netflix 0.921$\pm$ 0.220 1.861 $\pm$ 0.263 1.779 $\pm$ 0.241 1.888 $\pm$ 0.254 1.937 $\pm$ 0.258 1.933 $\pm$ 0.258 1.946 $\pm$ 0.264 1.919 $\pm$ 0.270* 1.910 $\pm$ 0.272*
Movielens 0.878$\pm$ 0.246 2.142 $\pm$ 0.114 2.080 $\pm$ 0.113 2.228 $\pm$ 0.046 2.232 $\pm$ 0.044 2.230 $\pm$ 0.046 2.235 $\pm$ 0.043 2.234 $\pm$ 0.046* 2.233 $\pm$ 0.050*
Yahoo 0.231 $\pm$ 0.248 1.684 $\pm$ 0.465 1.793 $\pm$ 0.466 1.793 $\pm$ 0.466 1.793 $\pm$ 0.466 1.793 $\pm$ 0.466 1.795 $\pm$ 0.466 1.729 $\pm$ 0.495* 1.715 $\pm$ 0.496*
LETOR 0.208 ± 0.106 0.440 ± 0.190 0.564 ± 0.184 0.577 ± 0.191 0.570 ± 0.191 0.557 ± 0.181 0.592 ± 0.195 0.582 ± 0.197 0.588 ± 0.197
LTRC 0.359 ± 0.116 0.597 ± 0.151 0.640 ± 0.134 0.662 ± 0.146 0.644 ± 0.145 0.642 ± 0.138 0.670 ± 0.144 0.667 ± 0.145 0.668 ± 0.145
[…]
Metrics Datasets Random EXPLORE DUM MSD MMR DPP B2I B3I B4I
[…]
KuaiRec 0.347 ± 0.049 0.483 ± 0.048 0.590 ± 0.048 0.459 ± 0.028 0.465 ± 0.030 0.472 ± 0.032 0.409 ± 0.072 0.477 ± 0.031* 0.480 ± 0.032*
Netflix 0.413 ± 0.065 0.575 ± 0.055 0.604 ± 0.044 0.596 ± 0.044 0.600 ± 0.044 0.600 ± 0.044 0.597 ± 0.046 0.600 ± 0.044* 0.600 ± 0.044*
Movielens 0.394 ± 0.080 0.620 ± 0.026 0.635 ± 0.005 0.635 ± 0.005 0.635 ± 0.005 0.635 ± 0.005 0.635 ± 0.005 0.635 ± 0.005* 0.635 ± 0.005*
Yahoo 0.360 ± 0.094 0.550 ± 0.100 0.572 ± 0.086 0.572 ± 0.086 0.572 ± 0.086 0.572 ± 0.086 0.565 ± 0.092 0.577 ± 0.085* 0.577 ± 0.085*
LETOR 0.274 ± 0.082 0.330±0.104 0.327 ± 0.104 0.416 ± 0.123 0.400 ± 0.115 0.406 ± 0.117 0.326 ± 0.104 0.390 ± 0.108 0.393 ± 0.110
LTRC 0.323 ± 0.068 0.373±0.082 0.378 ± 0.078 0.443 ± 0.084 0.447 ± 0.088 0.440 ± 0.081 0.378 ± 0.078 0.402 ± 0.080 0.407 ± 0.080
[…]
Netflix 0.028 ± 0.068 0.132±0.151 0.133 ± 0.171 0.152 ± 0.164 0.148 ± 0.164 0.149 ± 0.163 0.143 ± 0.161 0.139 ± 0.162* 0.139 ± 0.165*
Movielens 0.043 ± 0.091 0.141±0.153 0.136 ± 0.159 0.115 ± 0.144 0.115 ± 0.144 0.115 ± 0.144 0.115 ± 0.144 0.115 ± 0.144* 0.115 ± 0.144*
Yahoo 0.040 ± 0.079 0.275±0.145 0.309 ± 0.140 0.303 ± 0.136 0.303 ± 0.136 0.303 ± 0.136 0.325 ± 0.143 0.264 ± 0.143* 0.259 ± 0.145*

### v3

@@ 15    $\pi(2{\tau}-1)\leftarrow v_{2{\tau}-1}$, $\pi(2\tau)\leftarrow u_{2{\tau}-1}$
16 if $d\left(v_{2i-1},\pi(2i+1)\right)\geq d\left(u_{2i-1},\pi(2i+1)\right)$ then
@@ 6.2. The Greedy-matching Algorithm
In this section, we examine the case where $p$ is close to 1, but not a constant, specifically, when $\lim_{n\rightarrow\infty}p=1$ and $\lim_{n\rightarrow\infty}p^{n}=0$. This suggests that as the number of items increases, the user accepts more items, but eventually quits before examining all of them. For instance, setting $p=1-\frac{1}{\log\log n}$ implies that, in expectation, the user accepts $\log\log n$ items before quitting. For simplicity, we assume $p=1-\frac{1}{t_{n}}$, where $t_{n}$ is an increasing function of $n$.
In this scenario, applying $\small\textsc{B}\tau\textsc{I}$ algorithm to find best-$\tau$ items becomes computationally infeasible. As Theorem 1 shows, achieving a constant approximation requires setting $\tau=\Omega(t_{n})$. To address this, we introduce a polynomial-time greedy-matching (GM) algorithm.
@@
The next step is to extend $M$ to a Hamiltonian path. Notice that the (undirected) edge $(u_{2j-1},v_{2j-1})$ is always the (2j1)(2j-1)-th edge on the path, and the algorithm needs to decide which item comes first in the path. To make this decision, the algorithm utilizes the triangle inequality.
@@ 7. Non-uniform Continuation Probabilities
In this section, we discuss a more general case when the continuation probabilities $p_{i}$ are non-uniform. We assume $p_{i}\in[a,b]$, for all $i\in U$, where $0<a\leq b<1$. We start by adapting the $\small\textsc{B}\tau\textsc{I}$ algorithm, introduced in Section 6, to this setting. Given the computational complexity of the $\small\textsc{B}\tau\textsc{I}$ algorithm, we also propose an efficient greedy algorithm, whose performance guarantee is an implication of the $\small\textsc{B}\tau\textsc{I}$ algorithm by setting $\tau=2$.
@@ 7.1. The best-$\tau$ items Algorithm
Recall that the ordered Hamiltonian-path is defined as $\mathcal{H}(O)=\sum_{i=1}^{n-1}W_{O_{i}}d\left(\pi(i),\pi(i+1)\right)$, where $W_{O_{i}}=\sum_{j=i+1}^{n}p_{O_{j}}$. When the continuation probabilities $p_{i}$ are uniformly equal to $p$, we observe that the value of $W_{O_{i}}$ is dominated by $\nicefrac{{p^{i+1}}}{{(1-p)}}$, hereby, we can use $\nicefrac{{p^{i+1}}}{{(1-p)}}$ to approximate $W_{O_{i}}$. However, when $p_{i}$ are not uniform, it is hard to find a simple and closed-form equation to approximate $W_{O_{i}}$, which is convenient for our analysis.
@@
$\tilde{\mathcal{H}}(O_{\tau})=\sum\nolimits_{i=1}^{\tau-1}\sum\nolimits_{j=i+1}^{\tau}p_{O_{j}}d\left(\pi(i),\pi(i+1)\right),$
and we find the best-$\tau$ items by maximizing $\tilde{\mathcal{H}}(O_{\tau})$. We redefine
@@
The $\small\textsc{B}\tau\textsc{I}$ algorithm (Algorithm 1) has the following quality guarantee.
@@
Let $p_{i}\in[a,b]$ for all $i\in U$, where $0<a\leq b<1$. Let $\tau$ be an integer such that $2\leq\tau\leq n$. The $\small\textsc{B}\tau\textsc{I}$ algorithm provides a $\frac{a^{2}(1-b)(1-b^{\tau-1})}{a^{2}+(\tau-1)b^{\tau+1}}$-approximation guarantee for MaxOHP.
@@ 7.2. The Greedy Algorithm
Since the $\small\textsc{B}\tau\textsc{I}$ algorithm is impractical when $\tau$ is large, we propose a greedy algorithm, presented in Algorithm 3. Its approximation ratio is obtained by setting $\tau=2$ in Theorem 2.
@@
8. Directly Optimizing the MaxSSD Problem
In this section, we optimize $\mathcal{S}_{+}$ directly. We again consider two settings: when the continuation probability is uniform, and when it is non-uniform. For each setting, we propose a greedy algorithm and establish theoretical guarantees.
8.1. Uniform Continuation Probabilities
According to Equation 3.5, we can represent the sequential sum diversity of a sequence $O$, $\mathcal{S}_{+}(O)$, as follow:
(8.1) $\mathcal{S}_{+}(O)$ $=\sum_{k=1}^{n-1}\prod_{t=1}^{k}p_{\pi(t)}(1-p_{\pi(k+1)})\left[\sum\nolimits_{i,j\in O_{k}}d\left(i,j\right)\right]+\prod_{t=1}^{n}p_{\pi(t)}\left[\sum\nolimits_{i,j\in O}d\left(i,j\right)\right]$
Let $\operatorname{Div}(A)=\sum\nolimits_{i,j\in A}d\left(i,j\right)$ denote the max-sum diversification measure that has been well studied in the literature (Ravi et al., 1994; Gollapudi and Sharma, 2009; Borodin et al., 2012). When the item continuation probability $p$ is uniform across all items, Equation 8.1 is equivalent as follows.
(8.2) $\mathcal{S}_{+}(O)$ $=\sum_{k=1}^{n}\mathrm{Pr}(A=O_{k})\operatorname{Div}(O_{k})=\sum_{k=1}^{n-1}p^{k}(1-p)\operatorname{Div}(O_{k})+p^{n}\operatorname{Div}(O)$
We observe that: (1) when all items have the same continuation probability $p$, the coefficient $\mathrm{Pr}(A=O_{k})$ for the term $\operatorname{Div}(O_{k})$ is identical for all possible orderings $O$, and (2) $\mathrm{Pr}(A=O_{k})$ is non-increasing in $k\in\{1,\ldots,n-1\}$, i.e., top-rank items in the ordering $O$ contributes larger than low-rank items to the $\mathcal{S}_{+}$ value. Thus, to maximize $\mathcal{S}_{+}$, we can ignore the effect of $p$ and construct an ordering that greedily optimizes $\operatorname{Div}(O_{k})$ for all $k\in\{2,\cdots,n\}$. We show this method in Algorithm 4.
Input: Finite set $U$, probability $p$, diversity function $\operatorname{Div}$
Output: An ordered sequence of $U$
1 $S\leftarrow\emptyset$
2 while $|S|<n$ do
3 $u\leftarrow\operatorname*{arg\,max}_{u\in U\setminus S}\operatorname{Div}(S)$
4 $S\leftarrow S+u$
5
6return $S$
Algorithm 4 Greedy Algorithm for MaxSSD with uniform $p$ (Borodin et al., 2012)
Theorem 1.
Assume that $p_{i}=p$, for all $i\in U$. Algorithm 4 provides a $\nicefrac{{1}}{{2}}$-approximation guarantee for MaxSSD.
Proof.
Let $S^{*}_{k}=\operatorname*{arg\,max}_{S\subseteq U,|S|\leq k}\operatorname{Div}(S)$. According to Borodin et al. (2012), the classical greedy algorithm achieves a $\nicefrac{{1}}{{2}}$-approximation for the problem $\max_{S\subseteq U,|S|\leq k}\operatorname{Div}(S)$. Thus, let $S$ denote the solution returned by Algorithm 4, and let $S_{k}$ denote the obtained sequence of length $k$ after the $k$-th iteration. Then $\operatorname{Div}(S_{k})\geq\nicefrac{{1}}{{2}}\operatorname{Div}(S^{*}_{k})$ for all $k\in\{1,\ldots,n\}$.
Recall that $O^{*}$ denote the optimal sequence for $\mathcal{S}_{+}$: $O^{*}=\operatorname*{arg\,max}_{O=\pi(U)}\mathcal{S}_{+}(O)$. It holds that
(8.3) $\mathcal{S}_{+}(S)$ $=\sum_{k=1}^{n}\mathrm{Pr}(A=S_{k})\operatorname{Div}(S_{k})\stackrel{{\scriptstyle(a)}}{{\geq}}\frac{1}{2}\sum_{k=1}^{n}\mathrm{Pr}(A=S_{k})\operatorname{Div}(O^{*}_{k})$
$\stackrel{{\scriptstyle(b)}}{{=}}\frac{1}{2}\sum_{k=1}^{n}\mathrm{Pr}(A=O^{*}_{k})\operatorname{Div}(O^{*}_{k})\stackrel{{\scriptstyle(c)}}{{=}}\frac{1}{2}\mathcal{S}_{+}(O^{*})$
where inequality (a) holds because $\operatorname{Div}(S_{k})\geq\frac{1}{2}\operatorname{Div}(S^{*}_{k})\geq\frac{1}{2}\operatorname{Div}(O^{*}_{k})$ for all $k\in\{1,\ldots,n\}$. Equality (b) holds because the continuation probabilities are equal for all items, and (c) holds by the definition of $\mathcal{S}_{+}$.
8.2. Non-uniform Continuation Probabilities
When the continuation probabilities are non-uniform, the greedy algorithm in Algorithm 4 can perform poorly. Consider a worst-case scenario where the first two items selected by the greedy algorithm have continuation probabilities equal to 0, then the expected sequential diversity $\mathcal{S}_{+}$ is 0. In this case, we must account for the probabilities when constructing the ranking. To this end, we propose $\small\textsc{SB}\tau\textsc{I}$, which adapts the $\small\textsc{B}\tau\textsc{I}$ algorithm by selecting the best sequence of $\tau$ items that maximizes the sum of the first $\tau$ terms of $\mathcal{S}_{+}$, rather than selecting the best $\tau$ items that maximize $\tilde{\mathcal{H}}(O_{\tau})$. We denote this value as $\mathcal{S}_{+}(O_{\tau})$, which is defined as follows:
(8.4) $\mathcal{S}_{+}(O_{\tau})=\sum\nolimits_{i=1}^{\tau-1}p_{O_{i+1}}d\left(\pi(i+1),O_{i}\right).$
Theorem 2.
Let $p_{i}\in[a,b]$ for all $i\in U$, where $0<a\leq b<1$. Let $\tau$ be an integer such that $2\leq\tau\leq n$. The $\small\textsc{SB}\tau\textsc{I}$ algorithm, which essentially runs $\small\textsc{B}\tau\textsc{I}$ algorithm by setting $O_{\tau}=\operatorname*{arg\,max}_{O_{\tau}\sqsubseteq\pi(U)}\mathcal{S}_{+}(O_{\tau})$, provides a $\frac{a^{\tau}(1-b)(1-b^{\tau-1})}{a^{\tau}(1-b)+b^{\tau+1}}$-approximation guarantee for MaxSSD.
Proof.
Let $O^{*}$ denote the optimal ranking for the MaxSSD problem, and let $\pi^{*}$ denote the corresponding optimal permutation function. We partition the n1n-1 terms of $\mathcal{S}_{+}(O^{*})$ into $T=\lceil\frac{n}{\tau-1}\rceil$ chunks, with each of the first T1T-1 chunks containing $\tau-1$ terms, and denote the sum of the $j$-th chunk as $L_{j}$, which is calculated as follows:
$L_{j}=\sum_{i=(j-1)(\tau-1)+1}^{j(\tau-1)}\prod_{t=1}^{i+1}p_{\pi^{*}(t)}d\left(\pi^{*}(i+1),O^{*}_{i}\right)$
where we set $d\left(\pi^{*}(i+1),O^{*}(i)\right)=0$ for $i\geq n$, and it holds that $\mathcal{S}_{+}(O^{*})=\sum_{j=1}^{T}L_{j}$.
Let $L_{max}=\max_{O_{\tau-1}\sqsubseteq\pi(U)}\mathcal{S}_{+}(O_{\tau})$, then it holds that $L_{1}\leq L_{\max}$. Let $\ell=(j-1)(\tau-1)$, we can upper bound $L_{j}$ with $L_{\max}$ as follows
$L_{j}$ $=\prod_{k=1}^{\ell}p_{\pi^{*}(k)}\left[\sum_{i=\ell+1}^{j(\tau-1)}\prod_{t=\ell+1}^{i+1}p_{\pi^{*}(t)}d\left(\pi^{*}(i+1),O^{*}_{i}\right)\right]\leq b^{\ell}\left[\sum_{i=\ell+1}^{j(\tau-1)}\prod_{t=\ell+1}^{i+1}p_{\pi^{*}(t)}d\left(\pi^{*}(i+1),O^{*}_{i}\right)\right]$
$=b^{\ell}\Bigg[\sum_{i=\ell+1}^{j(\tau-1)}\prod_{t=\ell+1}^{i+1}p_{\pi^{*}(t)}d\left(\pi^{*}(i+1),\Big\{\pi^{*}(\ell+1),\cdots,\pi^{*}(i)\Big\}\right)$
$\qquad+\sum_{i=\ell+1}^{j(\tau-1)}\prod_{t=\ell+1}^{i+1}p_{\pi^{*}(t)}d\left(\pi^{*}(i+1),\Big\{\pi^{*}(1),\cdots,\pi^{*}(\ell)\Big\}\right)\Bigg]$
$\stackrel{{\scriptstyle(a)}}{{\leq}}b^{\ell}L_{\max}+\sum_{i=\ell+1}^{j(\tau-1)}b^{i+1}d\left(\pi^{*}(i+1),\Big\{\pi^{*}(1),\cdots,\pi^{*}(\ell)\Big\}\right)\stackrel{{\scriptstyle(b)}}{{\leq}}b^{\ell}L_{\max}+\frac{(j-1)b^{\ell}L_{\max}}{a^{\tau}}\sum_{i=2}^{\tau}b^{i}$
where inequality (a) holds by the definition of $L_{\max}$, and inequality (b) holds because for any $\tau$ items $\{x_{1},\ldots,x_{\tau}\}$, we have $d(x_{\tau},\{x_{1},\ldots,x_{\tau-1}\})\leq\sum_{i\neq j\leq\tau}d(x_{i},x_{j})=d(x_{1},x_{2})+d(x_{3},\{x_{1},x_{2}\})+\cdots+d(x_{\tau},\{x_{1},\ldots,x_{\tau-1}\})\leq\frac{L_{\max}}{a^{\tau}}$, and thus
$d\left(\pi^{*}(i+1),\Big\{\pi^{*}(1),\cdots,\pi^{*}((j-1)(\tau-1))\Big\}\right)$
$=\sum_{t=1}^{j-1}d(\pi^{*}(i+1),\Big\{\pi^{*}((t-1)(\tau-1)+1),\cdots,\pi^{*}(t(\tau-1))\Big\})\leq\frac{(j-1)L_{\max}}{a^{\tau}}$
Summing over all chunks, we obtain
$\mathcal{S}_{+}(O^{*})$ $=\sum_{j=1}^{T}L_{j}\leq\sum_{t=0}^{T-1}b^{(\tau-1)t}L_{\max}+(\sum_{t=2}^{\tau}b^{t})(\sum_{t=1}^{T-1}tb^{t(\tau-1)})\frac{L_{\max}}{a^{\tau}}$
$\stackrel{{\scriptstyle(a)}}{{=}}L_{\max}\left[\frac{1-b^{T(\tau-1)}}{1-b^{\tau-1}}+\frac{b^{2}(1-b^{\tau-1})}{a^{\tau}(1-b)}\cdot\frac{b^{\tau-1}\left(1-Tb^{(T-1)(\tau-1)}+(T-1)b^{T(\tau-1)}\right)}{(1-b^{\tau-1})^{2}}\right]$
$\leq L_{\max}\left[\frac{1}{1-b^{\tau-1}}+\frac{b^{2}(1-b^{\tau-1})}{a^{\tau}(1-b)}\cdot\frac{b^{\tau-1}}{(1-b^{\tau-1})^{2}}\right]=L_{\max}\frac{a^{\tau}(1-b)+b^{\tau+1}}{a^{\tau}(1-b)(1-b^{\tau-1})}$
where inequality (a) holds because
$\sum_{t=1}^{T-1}t\cdot b^{t}=\frac{b\left(1-T\cdot b^{(T-1)}+(T-1)b^{T}\right)}{(1-b)^{2}}$
is a standard arithmetico-geometric series summation, as documented in Equation 2.26 of Graham (1994).
Let $O$ denote the output of the $\small\textsc{B}\tau\textsc{I}$ algorithm with $O_{\tau}=\operatorname*{arg\,max}_{O_{\tau}\sqsubseteq\pi(U)}\mathcal{S}_{+}(O_{\tau})$. It holds that
$\frac{\mathcal{S}_{+}(O)}{\mathcal{S}_{+}(O^{*})}\geq\frac{L_{\max}}{L_{\max}\frac{a^{\tau}(1-b)+b^{\tau+1}}{a^{\tau}(1-b)(1-b^{\tau-1})}}\geq\frac{a^{\tau}(1-b)(1-b^{\tau-1})}{a^{\tau}(1-b)+b^{\tau+1}}$
8.3. $\small\textsc{B}\tau\textsc{I}$ and $\small\textsc{SB}\tau\textsc{I}$ Approximation Ratio Analysis
Let $A$ denote the approximation ratio of the $\small\textsc{SB}\tau\textsc{I}$ algorithm and $B$ denote the approximation ratio of the $\small\textsc{B}\tau\textsc{I}$ algorithm, where $A=\frac{a^{\tau}(1-b)(1-b^{\tau-1})}{a^{\tau}(1-b)+b^{\tau+1}}$ and $B=\frac{a^{3}(1-b)^{2}(1-b^{\tau-1})}{2b(1-a)(a^{2}+(\tau-1)b^{\tau+1})}.$ We observe that both approximation ratios $A$ and $B$ exhibit consistent monotonicity properties with respect to the parameters $a$ and $b$. Specifically, for fixed $b$, both $A$ and $B$ are monotonically increasing in $a$. Conversely, for fixed $a$, both $A$ and $B$ are monotonically decreasing in $b$. When both parameters increase simultaneously along the feasible region defined by $a\leq b$, the decreasing effect with respect to $b$ dominates, leading to an overall decrease in both approximation ratios. This dominance can be attributed to the higher-order dependence on $b$ in the denominators of both expressions.
To compare the performance of these two algorithms, we numerically compare their approximation ratios by visualizing ABA-B. We show the plots in Figure 1. For small $\tau$ (3, 4, 5), $A>B$ holds in the majority of the valid region (100.0 %, 99.7%, and 84.1%, respectively). As $\tau$ increases (5, 6, 7), the region where $A<B$ expands, with $A>B$ holding in only 65.5%, 52.1%, and 42.7% of the valid region, respectively. This suggests that $\small\textsc{SB}\tau\textsc{I}$ theoretically outperforms $\small\textsc{B}\tau\textsc{I}$ when $\tau$ is small, while $\small\textsc{B}\tau\textsc{I}$ becomes more competitive as $\tau$ increases. Consequently, neither algorithm consistently dominates the other, indicating that both approaches remain theoretically valid depending on the choice of $\tau$.
Figure 1. Heatmap of ABA-B for different values of $\tau$, where A is the approximation ratio of the $\small\textsc{SB}\tau\textsc{I}$ algorithm, and B the approximation ratio of the $\small\textsc{B}\tau\textsc{I}$ algorithm and $0<a\leq b<1$. Green regions indicate $A>B$ and red regions indicate $A<B$. The black contour line marks where $A=B$, and the dashed diagonal line represents the boundary $a=b$. The region below the dashed line is invalid as it violates the constraint $a\leq b$
@@ 9. Experimental Evaluation
All experiments are conducted on a dual-socket server, each socket housing an Intel Xeon Gold 6326 CPU at 2.90GHz, with 128 cores and 256 threads. Our implementation and the datasets are publicly available. [^1] https://github.com/HongLWang/Sequential-diversification-with-provable-guarantees
@@ 9.1. Datasets
We evaluate our methods on seven public datasets: five commonly used in recommender systems and two, LETOR and LTRC, from the information retrieval field. We refer to these datasets as recommendation datasets and information retrieval datasets, respectively. Table 1 provides a summary of dataset statistics. Due to space constraints, the dataset URLs are provided in the full paper.
@@
Table 1. Dataset statistics of the $5$ recommendation datasets and $2$ information retrieval datasets. $\mathcal{U}$ and $\mathcal{I}$ represent the set of users (or queries) and items (or documents), respectively. $\mathcal{R}$ denotes the number of ratings (or relevance scores), $\mathrm{avg}(\mathcal{R})$ denotes the average number of ratings per user (or documents per query), and $\mathrm{avg}(dist)$ denotes the average item-item (or document-document) distance.
Datasets $|\mathcal{U}|$ $|\mathcal{I}|$ $\mathcal{R}$ $\mathrm{avg}(\mathcal{R})$ $\mathrm{avg}(dist)$